OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

jimthompson5910 (jim_thompson5910):

Hint: multiply everything in the second equation by 2 to get 3x - 6y = 12 -4x + 6y = 12 then the equations add to get 3x - 6y = 12 -4x + 6y = 12 -------------- -1x+0y = 24 so -x = 24 ---> x = ??? then you use this to find y

OpenStudy (anonymous):

x= -24?

jimthompson5910 (jim_thompson5910):

now that you know x, you can find y

jimthompson5910 (jim_thompson5910):

so pick any equation, say the first one, then plug in x = -24 3x - 6y = 12 3(-24) - 6y = 12 -72 - 6y = 12 ... ... ... y = ??

OpenStudy (anonymous):

-14?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw